|
ID |
Date |
Author |
Subject |
Text |
 |
|
505
|
Sat Feb 10 17:05:13 2024 |
TD | Run 15 - background |
17.55 DAQ continues G22-86/R15_11
alpha background
|
20x |
|
504
|
Sat Feb 10 10:00:44 2024 |
AR RSS JM CB | Run 13, 14 | System checks: all ok
Attatchment 1: Rates ok
Attatchment 2: Temperatures ok |
23x |
|
503
|
Fri Feb 9 17:04:39 2024 |
JM, AR, CB, NH | Run12 - Background | |
|
|
502
|
Fri Feb 9 11:58:04 2024 |
TD + CB, JM, AR, NH, AB | Run10 contd, Run11 |
12.46 DAQ continues G22-88/R10_14
|
27x |
|
501
|
Fri Feb 9 07:57:38 2024 |
AR, CB, AB, JM | Run10 - First cool beam on target | 8:58
System wide checks okay, except
|
27x |
|
500
|
Thu Feb 8 17:11:50 2024 |
CB, EH | Pulser walkthrough |
17:30 File G22-88/R8
Pulser walkthrough
EGG
Ortec 448 Pulser
Amplitide 90,000-10,000
@ 10,000 step
Normalise 5.0
Attenuators |
 |
|
499
|
Thu Feb 8 14:59:59 2024 |
JM, TD | Change p+n FEE64 settings | ASIC settings for the pn FEEs has been reset
to the default aida settings. This was done
to combat the observed dropping of several |
|
|
498
|
Thu Feb 8 11:22:28 2024 |
JM | Shifters instruction 2024 | EMERGENCY SHUTDOWNS
1. CARME motors: Q key
on keyboard while on workspace 1 of appc218 |
|
|
497
|
Wed Feb 7 13:50:20 2024 |
Elena Hanu, Carlo Bruno, Michael Lestinsky | Afternoon shift | 14:50
Temperature, Statistics fine (see |
  |
|
496
|
Wed Feb 7 10:09:14 2024 |
CB, JG, EM | Morning Shift | 11:00
Changed slow comp threshold to
0x32 (500 keV). Checkloaded. |
   |
|
495
|
Tue Feb 6 09:39:54 2024 |
CB | Detector bias tests | 10:30 Found ADC#4 on aida 1 and 4 are suddenly
noisy. See attach 1-2. Checkloaded multiple
times - no effect. |
13x |
|
494
|
Mon Feb 5 20:15:31 2024 |
CB, TD | Detector move in/out signals | 17.43 setup signals from detector movement
controller.
|
|
|
493
|
Mon Feb 5 19:43:52 2024 |
CB, TD, ML | BuTIS reference for VETAR2 |
16.40 CB/ML install fibre optic cable from
patch panel beneath dipole downstream of |
  |
|
492
|
Mon Feb 5 12:48:50 2024 |
CB TD | Pulser walkthough - Run4 |
13.41 DAQ start - file G22-88/R4
|
 |
|
491
|
Mon Feb 5 05:45:26 2024 |
CB TD | Run3 - end | 06.40 DAQ continues file G22-88/R3_42
DSSSD bias OK - attachment 1
|
47x |
|
490
|
Sun Feb 4 08:40:04 2024 |
CB TD | Run3 - contd | 09.30 DAQ continues file G22-88/R3_17
DSSSD bias OK - attachment 1
|
22x |
|
489
|
Sat Feb 3 15:05:10 2024 |
CB, TD | [How To] Restart Anydesk | 1. Check Proxy server is correct
localhost : 8080
incoming connections 7070 |
|
|
488
|
Sat Feb 3 00:13:03 2024 |
CB TD | Run1 end - R3 start |
01.07 DAQ continues file G22-88/R1_46
background alpha
|
50x |
|
487
|
Fri Feb 2 09:58:58 2024 |
TD | TapeServer not writng data to disk - plus fix | DAQ and NewMerger appear to be cofigured correctly.
Can allocate device, create volume and open |
7x |
|
486
|
Fri Feb 2 08:59:36 2024 |
JM, TD, CB | Run1 start - pulser widths spreadsheet | Current pulser widths excel spreadsheet attached
09.58 DSSSD bias OK - attachment 1
|
20x |